The driver of the stationary car can hear the frequency of the sound either directly from the alarm of the car or from a bridge ahead.

(a.) Name the phenomenon that permits the driver of the stationary car to hear the sound from the bridge.

 (b.) At what frequencies does the driver of the stationary car hear the sounds?

 (c.) The stationary car starts moving towards the bridge at a speed of 6ms~1. At what frequencies does the driver now hears the sounds? Explain why he hears beats and calculate the beat frequency


1
Expert's answer
2021-03-23T08:19:47-0400

(a.) The phenomenon that permits the driver of the stationary car to hear the sound from the bridge is called echo.

(b.) The driver of the stationary car hears the sounds at frequencies between 20 Hz to 20 kHz.

(c.) When the stationary car starts moving towards the bridge at a speed of 6m/s, the driver hears the frequencies of


"f_l=f_{l0}(1+v/c)=20(1+6/343)=20.3\text{ Hz},\\
f_u=f_{u0}(1+v/c)=20(1+6/343)=20.3\text{ kHz}."